OpenStudy (anonymous):

josh and mae work at a concession stand. thy each earn $8 per hour. josh worked 3 hours more than mae. if josh and mae earned a total of $120 , how many hours did josh work ???

OpenStudy (pfenn1):

Let x = number of hours Josh worked and y = number of hours Mae worked.\[x=3+y\]\[8(x+y)=120\]

OpenStudy (anonymous):

alright so both workers make 8 dollars per hour allowing E equal earnings and x equal hours you get the equation E=8x+8y, [each person makes 8 so if two are working , you're earning 16 dollars per hour now same it gives you the earnings for the second equation 120 but it also says that one of them worked 3 more hours. if x=hours that one worked, y=x+3 must be how much the other worked using this you get 120=8x+8y 120=8x+8(x+3)

OpenStudy (anonymous):

once you find x you just need to find y by plugging into the other equation